# Quantum correlations under decoherence

### v1

@@ V Results and discussion
Figure 4: Time of entanglement death, $\gamma\tau$, versus $\bar{n}$ (left panel) and concurrence (right panel) versus $\gamma t$ for $\bar{n}=0$ (black), $\bar{n}=0.2$ (red), $\bar{n}=0.4$ (green) and $\bar{n}=0.6$ (blue) for $\omega=1$ and $\Delta=J=0$ i.e., qubits solely interact with the common environment not with each other. System is initially described by the mixture (91) with $w=\frac{1}{2}$.
Next, we illustrate in figure 5 dynamics of quantum correlations for the Werner state (62) for different values of $p$. Initial values of quantum correlations are $\mathcal{C}(0)=\max\{0,|p|-\frac{1-p}{2}\}$, $\mathcal{L_{N}}(0)=\log_{2}\left(\frac{1+p}{2}+\frac{1}{2}\left|\frac{1-p}{2}-|p|\right|+\frac{1}{2}\left|\frac{1-p}{2}+|p|\right|\right)$, $C_{cc}(0)=|p|$ and $\mathcal{U}(0)=1-\frac{1}{2}[1-p+\sqrt{(1-p)(1+3p)}]$ from which one finds the concurrence is non-zero only for $\frac{1}{3}<p\leq 1$ implying, as is well-known, that in this region the Werner state (62) is entangled. While QE, quantified by both concurrence and log negativity, is initially zero for values $p=-\frac{1}{3}$ and $p=0$ after a while it is born and finally reaches the steady value $0.2999$ and $0.3784$, being independent of $p$, respectively for concurrence and log negativity for $\gamma=0.1$, $\omega=1$ and $\Delta=0.5$. Using Eq. (74) one obtains $\mathcal{C}(\infty)=2\max\left\{0,\frac{|\Delta|\sqrt{4\omega^{2}+\gamma^{2}}-\Delta^{2}}{4\Omega^{2}+\gamma^{2}}\right\}$ for the steady state value of quantum concurrence which is non-zero only for $|\Delta|<\sqrt{4\omega^{2}+\gamma^{2}}$ As pointed out in the previous section, for $\gamma=0$, the Werner state (62) does not evolve in time meaning that all quantum correlations remain constant. From all these we see that if both conditions $\frac{1}{3}<p\leq 1$ and $|\Delta|<\sqrt{4\omega^{2}+\gamma^{2}}$ fulfil simultaneously then for $|p|-\frac{1-p}{2}<\frac{|\Delta|\sqrt{4\omega^{2}+\gamma^{2}}-\Delta^{2}}{4\Omega^{2}+\gamma^{2}}$, the steady state is more entangled than the initial state. For parameters $\gamma=0.1$, $\omega=1$ and $\Delta=0.5$ where $\mathcal{C}(\infty)=0.2999$, this condition fulfils for $\frac{1}{3}<p<0.5333$.
Figure 5: Evolution of the quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel), for $p=-\frac{1}{3}$ (black curves), $p=0$ (red curves), $p=\frac{1}{\sqrt{3}}$ (green curves), and $p=1$ (blue curves) for $\gamma=0.1$, $\omega=1$ and $\Delta=0.5$. Note that the Werner state is independent of the isotropic part of qubits’ interaction i.e., the coefficient $J$.
Figure 6: Evolution of the quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel) for different values of the anisotropy interaction coefficient: $\Delta=0$ (brown), $\Delta=0.2$ (magenta) $\Delta=0.3$ (cyan) and $\Delta=0.4$ (orange); and for $\omega=1$ and $\gamma=0.1$. System is initially described by the Werner state (62) with $p=0$ (completely random ensemble; $\rho_{0}=\frac{1}{4}\mathds{1}_{4}$).
After, we highlight the constructive role played by the environment in favour of quantum correlations; the so-called environment-induced or decoherence-induced quantum correlations. To this end, we consider our system is initially described by the Werner state (62) with $p=0$ i.e., $\rho_{0}=\frac{1}{4}\mathds{1}_{4}$; the completely random ensemble which contains no correlations. Then, evolution of quantum correlations is considered for different values of the interaction strength $\Delta$. Note that the evolved state and thus quantum correlations are all independent of the value of $J$. Figure 6 depicts variation of quantum correlations and CC with time for a given value of $\gamma$ but different values of interaction strength $\Delta$. For $\Delta=0$ all quantum correlations are always zero but for $\Delta\neq 0$ quantum correlations are generated as time elapses. From this point we learn that (i) environment plays a constructive role, in contrast to its customary destructive role, inducing quantum correlations (ii) in addition to the interaction of qubits with the environment, the anisotropic part of their own interaction with each other is necessary for creation of quantum correlations. As top plots of this figure show, generation of QE from an initially separable state, the so-called environment-induced entanglement, requires both interaction between qubits and also their own interaction with the environment; no one is sufficient on its own. This constructive role played by the environment in favour of entanglement, in contrast to its customary destructive role, has already been noticed. See for instance AnWaLu-PA-2007 ; Hoetala-PRA-2009 . Despite the apparent implication from this figure that quantum correlations increase with the strength of interaction $\Delta$, it is important to note that this is not always the case. Consider the steady state as a example. For our parameters, $\gamma=0.1$ and $\omega=1$, the steady state concurrence $\mathcal{C}(\infty)=2\max\left\{0,\frac{|\Delta|\sqrt{4\omega^{2}+\gamma^{2}}-\Delta^{2}}{4\Omega^{2}+\gamma^{2}}\right\}$ is nonzero only for $\Delta\in(-2.0025,0)\cup(0,2.0025)$ and is an increasing function of $\Delta$ only when $-2.0025<\Delta\leq-0.6188$ and $0<\Delta\leq 0.6188$.
Figure 7: Plots of the finite temperature, steady state quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel), in terms of the average excitation of the thermal bath, $\bar{n}$, for $\gamma=0.01$ (blue curves) and $\gamma=3$ (orange curves) for $\Delta=0.5$ and $\omega=1$. Note that X-shaped steady state is independent of both the initial state and the isotropic part of qubits’ interaction i.e., the coefficient $J$.
@@
Finally, in figure 7 we have plotted the finite temperature, steady state quantum correlations versus the average excitation of the thermal bath, $\bar{n}$, for two values of the emission rate, $\gamma=0.01$ (blue curves) and $\gamma=3$ (orange curves), for $\Delta=0.5$ and $\omega=1$. As one sees, a steady state level of all quantum correlations is always reached for zero or finite, low temperatures. For nonzero temperatures all quantum correlations are found to vanish at a finite temperature. This means none of the quantum correlations are robust against the temperature. However, CC which equals MIN is more robust than others. Emission rate $\gamma$ speeds up the decay of quantum correlations.
@@ VI Summary and conclusions
In this work, our model was a system of two localized qubits interacting through their spins, and also, an external magnetic field was applied. The interaction with the environment was introduced by a non-unitary term in the Lindblad equation. We solved this equation analytically at zero temperature for several X-shaped states and studied the evolution of different quantum correlations for different values of the emission rate $\gamma$. After that, it was time to bring the temperature into play and consider the influence of thermal decoherence on quantum correlations. Although the corresponding master equation could be solved analytically, but due to the length of the solution, we considered only the stable solution and examined the dependence of different quantum correlations on temperature in this case. Our main results can be summarized as follows: (i) MIN is equal to CC. (ii) The sudden death of quantum correlations beyond entanglement is not possible, maybe a not an unexpected result with respect to their definition. (iii) Even if the initial state lacks correlations, the interaction with the environment induces them. In other words, like entanglement, it is possible to induce other quantum correlations by the decoherence. (iv) In addition to the interaction of qubits with the common environment, their own interaction with each other is necessary to induce quantum correlations. (v) Although the revival and rebirth of entanglement happens after a finite period of time, the revival of other correlations is much faster. (vi) In the zero-temperature steady state, LQU captures less quantum correlations than the other quantifiers. (vii) Temperature speeds up ESD. (viii) Quantum correlations beyond entanglement are more resistant to the negative effect of temperature to destroy them.

### v2

@@ V Results and discussion
Figure 6: Time of entanglement death, $\gamma\tau$, versus $\bar{n}$ (left panel) and concurrence (right panel) versus $\gamma t$ for $\bar{n}=0$ (black), $\bar{n}=0.2$ (red), $\bar{n}=0.4$ (green) and $\bar{n}=0.6$ (blue) for $\Delta=J=0$ i.e., qubits solely interact with the common environment not with each other. System is initially described by the mixture (83) with $w=\frac{1}{2}$.
@@
Finally, in figure 7 we have plotted the finite temperature, steady state quantum correlations versus the average excitation of the thermal bath, $\bar{n}$, for two values of the emission rate, $\gamma=0.01\omega$ (magenta curves) and $\gamma=3\omega$ (orange curves), for $\Delta=0.5\omega$. As one sees, a steady state level of all quantum correlations is always reached for zero or finite, low temperatures. For nonzero temperatures, all quantum correlations are found to vanish at a finite temperature. This means none of the quantum correlations are robust against the temperature. However, CC which equals MIN is more robust than others. Emission rate $\gamma$ speeds up the decay of quantum correlations.
Figure 7: Plots of the finite temperature, steady state quantum correlations, concurrence (left top panel), log negativity (right top panel), LQU (left bottom panel) and CC (right bottom panel), in terms of the average excitation of the thermal bath, $\bar{n}$, for $\gamma=0.01\omega$ (blue curves) and $\gamma=3\omega$ (orange curves) for $\Delta=0.5\omega$. Note that X-shaped steady state is independent of both the initial state and the isotropic part of qubits’ interaction, i.e., the coefficient $J$.
@@ VI Summary and conclusions
In this work, our model was a system of two localized qubits interacting through their spins, and also, an external magnetic field was applied. The interaction with the environment was introduced by a non-unitary term in the Lindblad equation. We solved this equation analytically at zero temperature for several X-shaped states and studied the evolution of different quantum correlations for different values of the emission rate $\gamma$. After that, it was time to bring the temperature into play and consider the influence of thermal decoherence on quantum correlations. Although the corresponding master equation could be solved analytically, but due to the length of the solution, we considered only the stable solution and examined the dependence of different quantum correlations on temperature in this case. Our main results can be summarized as follows: (i) MIN is equal to CC. (ii) The sudden death of quantum correlations beyond entanglement is not possible, maybe not an unexpected result with respect to their definition. (iii) Even if the initial state lacks correlations, the interaction with the environment induces them. In other words, like entanglement, it is possible to induce other quantum correlations by the decoherence. (iv) In addition to the interaction of qubits with the common environment, their own interaction with each other is necessary to induce quantum correlations. (v) Although the revival and rebirth of entanglement happens after a finite period of time, the revival of other correlations is much faster. (vi) In the zero-temperature steady state, LQU captures less quantum correlations than the other quantifiers. (vii) Temperature speeds up ESD. (viii) Quantum correlations beyond entanglement are more resistant to the negative effect of temperature to destroy them.